OpenStudy (anonymous):

Slope of 0, through the point (9,-7)

OpenStudy (luigi0210):

Use \(y-y_{1}=m(x-x_{1})\) where \(m=slope\)

OpenStudy (solomonzelman):

only vertical line has a slope of zero, a vertical line can be y=3 where the points on the line can have any x value but the y value will be always 3 (anything, 3 ) Which vertical line has a point (9,-7) ? (think, the y is supposed to be equal to -7 (always) ).

OpenStudy (ipwnbunnies):

Err, you mean horizontal line.

OpenStudy (solomonzelman):

Yeah, tnx for correcting me :)